OpenStudy (explainitlikeimfive):

Which of the following is the reciprocal of x2 – 2x – 8? A. (x + 2)(x – 4) B. 1 (x + 2)(x – 4) C. 1 (x – 8) D. 1 (x + 2)

OpenStudy (anonymous):

Given: \[x^2 – 2x – 8 \rightarrow \frac{ x^2 – 2x – 8 }{ 1 }\]

OpenStudy (anonymous):

1.) factor \(x^2 – 2x – 8 \)

OpenStudy (anonymous):

HINT: what two numbers when multiplied give you (-8) but when added together give you (-2)?

OpenStudy (anonymous):

@ExplainItLikeImFive

OpenStudy (explainitlikeimfive):

-4 and 2

OpenStudy (anonymous):

Good

OpenStudy (anonymous):

\[ x^2 – 2x – 8 = (x + 2)(x-4)\]

OpenStudy (anonymous):

Given: \[\frac{ x^2 – 2x – 8 }{ 1 }\] Reciprocal: \[\frac{ 1 }{ x^2 - 2x - 8 } \rightarrow \frac{ 1 }{ (x+2)(x-4) } \]

OpenStudy (anonymous):

Do you understand @ExplainItLikeImFive ?

OpenStudy (explainitlikeimfive):

Yes, you did a fantastic job at explaining it. Thanks

OpenStudy (anonymous):

Reciprocal is when you 'flip it'

OpenStudy (anonymous):

Thank you :) & Your Welcome :)